OpenStudy (anonymous):

Choose the correct simplification of (3xy6)2(y3)3. 9x2y21 6x2y21 9x2y14 6x2y14

OpenStudy (freckles):

is this what you mean: \[(3xy^6)^2(y^3)^3\]?

OpenStudy (anonymous):

9x^2y^21 6x^2y^21 9x^2y^14 6x^2y^14

OpenStudy (anonymous):

Yes. Sorry 1 just copy and pasted it.

OpenStudy (freckles):

apply this rule first: \[(a^r x^sy^t)^z=a^{r \cdot z}x^{s \cdot z}y^{t \cdot z}\]

OpenStudy (freckles):

\[(3^1 x^1y^6)^2=? \\ (y^3)^3=?\]

OpenStudy (freckles):

this rule says you are going to multiply the exponents inside by the outer exponent

OpenStudy (anonymous):

\[3x^2y^{12} \] and\[y^9\]

OpenStudy (freckles):

well you just forgot to do it to 1st power on the 3

OpenStudy (anonymous):

1 thought 3 stayed the same?

OpenStudy (freckles):

\[(3^1 x^1y^6)^2=3^{1 \cdot 2}x^{1 \cdot 2}y^{6 \cdot 2 }=3^{2}x^{2}y^{12} \\ (y^3)^3=y^{3 \cdot 3}=y^{9} \\ \text{ so now you have } 3^2 x^2 y^{12} \cdot y^9\] -- why would 3 stay the same when it is being squared?

OpenStudy (freckles):

\[3^2 \neq 3 \]

OpenStudy (freckles):

\[3^2=3(3)=?\]

OpenStudy (anonymous):

okay so the first one?

OpenStudy (freckles):

right since \[y^{12} y^9=y^{12+9}=y^{21}\]

OpenStudy (freckles):

just so you know x and y are numbers just like 3 is a number

OpenStudy (anonymous):

can you help with another?

OpenStudy (anonymous):

okay

OpenStudy (freckles):

sure

OpenStudy (anonymous):

Choose the correct simplification of \[\frac{ f ^{9} h ^{23} }{ f ^{3} h ^{17} }\]

OpenStudy (freckles):

use quotient rule for exponents here

OpenStudy (freckles):

\[\frac{u^{n}}{u^m}=u^{n-m}\]

OpenStudy (freckles):

\[\frac{f^9}{f^3} \cdot \frac{h^{23}}{h^{17}}=?\]

OpenStudy (anonymous):

\[f ^{6}h ^{6}\]

OpenStudy (freckles):

yep since 9-3=6 and 23-17=6

OpenStudy (anonymous):

awesome thanks so much!

OpenStudy (freckles):

np